If the volume of a gas is 19.4 L at 4.99 atm, what is the volume at 1580 mmHg


1
Expert's answer
2021-06-15T05:13:30-0400

"V_1=19.4\;L \\

P_1=4.99 \;atm \\

P_2=1580 \;mmHg = 2.078 \;atm \\

P_1V_1=P_2V_2 \\

V_2= \frac{P_1V_1}{P_2} \\

= \frac{4.99 \times 19.4}{2.078} = 46.58 \;L"

Answer: 46.58 L